# BioNeRF: Biologically Plausible Neural Radiance Fields for View Synthesis

### v2

@@ 6.1 General Evaluation
NeRF mildenhall2021nerf (ACM 21) $31.01$ $33.00$ $25.01$ $30.13$ $36.18$ $32.54$ $29.62$ $32.91$ $28.65$
TensoRF chen2022tensorf (ECCV’22) $33.14$ $35.76$ $26.01$ $33.99$ $37.41$ $\bm{36.46}$ $30.12$ $34.61$ $30.77$
Plenoxels fridovich2022plenoxels (CVPR’22) $31.71$ $33.98$ $25.35$ $31.83$ $36.43$ $34.10$ $29.14$ $33.26$ $29.62$
Point-NeRF xu2022point (CVPR’22) $33.31$ $35.40$ $\bm{26.06}$ $\bm{36.13}$ $37.30$ $35.04$ $29.61$ $35.95$ $30.97$
Ref-NeRF verbin2022ref (CVPR’22) $\bm{33.99}$ $\bm{35.83}$ $25.79$ $33.91$ $\bm{37.72}$ $36.25$ $\bm{35.41}$ $\bm{36.76}$ $30.28$
L2G-NeRF chen2023local (CVPR’23) $28.62$ $30.99$ $23.75$ $26.11$ $34.56$ $27.71$ $27.60$ $30.91$ $27.31$
Tetra-NeRF kulhanek2023tetra (ICCV’23) $32.53$ $35.05$ $25.01$ $33.31$ $36.16$ $34.75$ $29.30$ $35.49$ $\bm{31.13}$
@@
NeRF mildenhall2021nerf (ACM 21) $0.947$ $0.967$ $0.925$ $0.964$ $0.974$ $0.961$ $0.949$ $0.980$ $0.856$
TensoRF chen2022tensorf (ECCV’22) $0.963$ $0.985$ $0.937$ $0.982$ $0.982$ $0.983$ $0.952$ $0.988$ $0.895$
Plenoxels fridovich2022plenoxels (CVPR’22) $0.958$ $0.977$ $0.933$ $0.890$ $0.985$ $0.976$ $\bm{0.975}$ $0.980$ $0.949$
Point-NeRF xu2022point (CVPR’22) $0.978$ $\bm{0.991}$ $\bm{0.954}$ $\bm{0.993}$ $\bm{0.991}$ $\bm{0.988}$ $0.971$ $\bm{0.994}$ $0.942$
Ref-NeRF verbin2022ref (CVPR’22) $0.966$ $0.984$ $0.937$ $0.983$ $0.984$ $0.981$ $0.983$ $0.992$ $0.880$
L2G-NeRF chen2023local (CVPR’23) $0.930$ $0.950$ $0.900$ $0.930$ $0.970$ $0.910$ $0.930$ $0.970$ $0.850$
Tetra-NeRF kulhanek2023tetra (ICCV’23) $\bm{0.982}$ $0.990$ $0.947$ $0.989$ $0.989$ $0.987$ $0.968$ $0.993$ $\bm{0.994}$
@@
NeRF mildenhall2021nerf (ACM 21) $0.081$ $0.046$ $0.091$ $0.044$ $0.121$ $0.050$ $0.063$ $0.028$ $0.206$
TensoRF chen2022tensorf (ECCV’22) $0.027$ $\bm{0.010}$ $0.051$ $\bm{0.012}$ $0.013$ $\bm{0.007}$ $0.026$ $\bm{0.009}$ $0.085$
Plenoxels fridovich2022plenoxels (CVPR’22) $0.049$ $0.031$ $0.067$ $0.026$ $0.037$ $0.028$ $0.057$ $0.015$ $0.134$
Point-NeRF xu2022point (CVPR’22) $0.049$ $0.023$ $0.078$ $0.022$ $0.037$ $0.024$ $0.072$ $0.014$ $0.124$
Ref-NeRF verbin2022ref (CVPR’22) $0.038$ $0.017$ $0.059$ $0.019$ $0.022$ $0.018$ $0.022$ $0.007$ $0.139$
L2G-NeRF chen2023local (CVPR’23) $0.070$ $0.050$ $0.100$ $0.060$ $0.030$ $0.060$ $0.060$ $0.050$ $0.130$
Tetra-NeRF kulhanek2023tetra (ICCV’23) $0.041$ $0.016$ $0.073$ $0.023$ $0.027$ $0.022$ $0.056$ $0.011$ $0.103$
@@
6.2 Comparisons on Realistic Synthetic ${}^{\circ}$ dataset.
Table 2 presents results concerning the Realistic Synthetic ${}^{\circ}$ dataset. Regarding the PSNR measure, one can observe that Ref-NeRF obtained the most accurate average results, achieving higher PSNR values over half of the scenes, while Point-NeRF stood out in two of them, and TensoRF and Tetra-NeRF did well in Lego and Ship, respectively. Concerning the SSIM metric, even though Point-NeRF obtained the best results on six out of eight scenes, Tetra-NeRF achieved the best average accuracy overall due to an excellent performance over the Ship scene.
Apart from such results, BioNeRF asserted the average best results considering the LPIPS measure, obtaining the minimum value over four out of eight scenes, i.e., Drums, Hotdog, Materials, and Ship. Further, it virtually broke even with the best results considering the Chair scene, reaching an LPIPS of $0.011$ against $0.010$ from TensoRF. Such results are very positive since LPIPS is the measure that best corresponds to human perceptual judgments, thus implying better visual appearance. Figure 4 depicts some examples from BioNeRF view synthesis.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption drums hotdog materials ship
Figure 4: Ground-truth (top) and synthetic view (bottom) images generated by BioNeRF regarding four Realistic Synthetic ${}^{\circ}$ dataset’s scenes.
6.3 Comparisons on LLFF Real dataset.
@@
NeRF mildenhall2021nerf (ACM 21) $26.50$ $25.17$ $27.40$ $31.16$ $27.45$ $20.92$ $20.36$ $\bm{32.70}$ $26.80$
TensoRF chen2022tensorf (ECCV’22) $26.73$ $25.27$ $\bm{28.60}$ $31.36$ $\bm{28.14}$ $21.30$ $19.87$ $32.35$ $26.97$
Plenoxels fridovich2022plenoxels (CVPR’22) $26.29$ $25.46$ $27.83$ $31.09$ $27.58$ $21.41$ $20.24$ $30.22$ $26.48$
L2G-NeRF chen2023local (CVPR’23) $24.54$ $24.57$ $24.90$ $29.27$ $23.12$ $19.02$ $19.71$ $32.25$ $23.49$
BioNeRF (Ours) $\bm{27.01}$ $\bm{26.51}$ $27.89$ $\bm{32.34}$ $27.99$ $\bm{22.23}$ $\bm{20.80}$ $30.75$ $\bm{27.56}$
SSIM normal-$\uparrow$
@@
NeRF mildenhall2021nerf (ACM 21) $0.811$ $0.792$ $0.827$ $0.881$ $0.828$ $0.690$ $0.641$ $0.948$ $0.880$
TensoRF chen2022tensorf (ECCV’22) $0.839$ $0.814$ $0.871$ $0.897$ $0.877$ $0.752$ $0.649$ $0.952$ $0.900$
Plenoxels fridovich2022plenoxels (CVPR’22) $0.839$ $0.832$ $0.862$ $0.885$ $0.857$ $0.760$ $0.687$ $0.937$ $0.890$
L2G-NeRF chen2023local (CVPR’23) $0.750$ $0.750$ $0.740$ $0.840$ $0.740$ $0.560$ $0.610$ $0.950$ $0.800$
BioNeRF (Ours) $\bm{0.861}$ $\bm{0.837}$ $\bm{0.873}$ $\bm{0.914}$ $\bm{0.882}$ $\bm{0.796}$ $\bm{0.714}$ $\bm{0.956}$ $\bm{0.911}$
LPIPS normal-$\downarrow$
@@
NeRF mildenhall2021nerf (ACM 21) $0.250$ $0.280$ $0.219$ $0.171$ $0.268$ $0.316$ $0.321$ $0.178$ $0.249$
TensoRF chen2022tensorf (ECCV’22) $0.124$ $0.155$ $0.106$ $0.075$ $0.123$ $0.153$ $0.201$ $0.082$ $0.099$
Plenoxels fridovich2022plenoxels (CVPR’22) $0.210$ $0.224$ $0.179$ $0.180$ $0.231$ $0.198$ $0.242$ $0.192$ $0.238$
L2G-NeRF chen2023local (CVPR’23) $0.200$ $0.260$ $0.170$ $0.110$ $0.260$ $0.330$ $0.250$ $0.080$ $0.160$
@@
Figure 5 depicts some high-quality views generated from LLFF dataset scenes. Such samples confirm BioNeRF’s robustness in generating pleasant images for human perceptual judgments.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption fern trex horns flower
Figure 5: Ground-truth (top) and synthetic view (bottom) images generated by BioNeRF regarding four LLFF dataset’s scenes.
6.4 Neuronal Activity Analysis
This section briefly analyzes the neuronal activity concerning $M_{\Delta}$, $M_{c}$, $M^{\prime}_{\Delta}$, and $M^{\prime}_{c}$. Figure 6 depicts the average rate of neuronal activation per model and their respective standard deviation for both datasets considering the first $10,000$ iterations. In both cases, models $M_{\Delta}$ and $M_{c}$ present a higher neuronal activity in initial iterations, decreasing as time passes. Such a behavior can be explained by learning more generic features at the beginning of the process and refining them with more intrinsic and subtle features as training converges.
Regarding the $M^{\prime}_{\Delta}$ and $M^{\prime}_{c}$, there is an opposite behavior on both datasets, i.e., higher activity concerning $M^{\prime}_{\Delta}$ and lower activity concerning $M^{\prime}_{c}$ over Realistic Synthetic ${}^{\circ}$, and an inversed conduct considering LLFF. Such a behavior sheds light on the complexity of each dataset: Realistic Synthetic ${}^{\circ}$ comprises single, artificially modeled objects, consecutively demanding less effort to learn the density of the scenes but requiring more resources to estimate the reflected colors. On the other hand, LLFF comprises scenes composed of several layers of objects and backgrounds, imposing more efforts to learn the scenes densities while alleviating the color estimation burden since the images were captured in a smaller range of angles, resulting in a lighter spectrum of reflected lights.
Refer to caption Refer to caption
Figure 6: Average neuronal activity and standard deviation for neural models $M_{\Delta}$, $M_{c}$, $M^{\prime}_{\Delta}$, and $M^{\prime}_{c}$ over the (left) Realistic Synthetic ${}^{\circ}$ and (right) LLFF datasets.
@@ 7 Discussions, Conclusions, and Future Works
Future Works. Regarding future works, we aim to improve the model by introducing more elements from a biologically plausible perspective, like spiking neurons, implemented in SpikingNeRF yao2023spiking . Additionally, we strive to extend our model to infer in real time and learn from a few shots.
Acknowledgments
The authors are grateful to Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Brazil grants #429003/2018-8, #307066/2017-7 and #427968/2018-6, Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP), Brazil grants #2021/05516-1, #2013/07375-0, #2014/12236-1, #2023/10823-6, and #2019/07665-4.

### v3

@@ 5.2 Quantitative evaluation
NeRF [16] (ACM 21) $31.01$ $33.00$ $25.01$ $30.13$ $36.18$ $32.54$ $29.62$ $32.91$ $28.65$
[…]
Plenoxels [8] (CVPR’22) $31.71$ $33.98$ $25.35$ $31.83$ $36.43$ $34.10$ $29.14$ $33.26$ $29.62$
Point-NeRF [24] (CVPR’22) $33.31$ $35.40$ $\bm{26.06}$ $\bm{36.13}$ $37.30$ $35.04$ $29.61$ $35.95$ $30.97$
Ref-NeRF [22] (CVPR’22) $\bm{33.99}$ $\bm{35.83}$ $25.79$ $33.91$ $\bm{37.72}$ $36.25$ $\bm{35.41}$ $\bm{36.76}$ $30.28$
Mip-NeRF 360 [4] (CVPR’22) $30.34$ $34.01$ $24.36$ $26.66$ $36.44$ $33.20$ $27.91$ $31.50$ $28.66$
L2G-NeRF [7] (CVPR’23) $28.62$ $30.99$ $23.75$ $26.11$ $34.56$ $27.71$ $27.60$ $30.91$ $27.31$
Tetra-NeRF [14] (ICCV’23) $32.53$ $35.05$ $25.01$ $33.31$ $36.16$ $34.75$ $29.30$ $35.49$ $\bm{31.13}$
@@
NeRF [16] (ACM 21) $0.947$ $0.967$ $0.925$ $0.964$ $0.974$ $0.961$ $0.949$ $0.980$ $0.856$
TensoRF [6] (ECCV’22) $0.963$ $0.985$ $0.937$ $0.982$ $0.982$ $0.983$ $0.952$ $0.988$ $0.895$
[…]
Point-NeRF [24] (CVPR’22) $0.978$ $\bm{0.991}$ $\bm{0.954}$ $\bm{0.993}$ $\bm{0.991}$ $\bm{0.988}$ $0.971$ $\bm{0.994}$ $0.942$
Ref-NeRF [22] (CVPR’22) $0.966$ $0.984$ $0.937$ $0.983$ $0.984$ $0.981$ $\bm{0.983}$ $0.992$ $0.880$
Mip-NeRF 360 [4] (CVPR’22) $0.951$ $0.977$ $0.923$ $0.952$ $0.979$ $0.975$ $0.944$ $0.984$ $0.875$
L2G-NeRF [7] (CVPR’23) $0.930$ $0.950$ $0.900$ $0.930$ $0.970$ $0.910$ $0.930$ $0.970$ $0.850$
[…]
@@
NeRF [16] (ACM 21) $0.081$ $0.046$ $0.091$ $0.044$ $0.121$ $0.050$ $0.063$ $0.028$ $0.206$
[…]
Plenoxels [8] (CVPR’22) $0.049$ $0.031$ $0.067$ $0.026$ $0.037$ $0.028$ $0.057$ $0.015$ $0.134$
Point-NeRF [24] (CVPR’22) $0.049$ $0.023$ $0.078$ $0.022$ $0.037$ $0.024$ $0.072$ $0.014$ $0.124$
[…]
Mip-NeRF 360 [4] (CVPR’22) $0.060$ $0.032$ $0.083$ $0.048$ $0.039$ $0.028$ $0.067$ $0.021$ $0.164$
L2G-NeRF [7] (CVPR’23) $0.070$ $0.050$ $0.100$ $0.060$ $0.030$ $0.060$ $0.060$ $0.050$ $0.130$
[…]
@@
Apart from such results, BioNeRF asserted the average best results considering the LPIPS measure, obtaining the minimum value over four out of eight scenes, i.e., Drums, Hotdog, Materials, and Ship. Further, it virtually broke even with the best results considering the Chair scene, reaching an LPIPS of $0.011$ against $0.010$ from TensoRF. Such results are very positive since LPIPS is the measure that best corresponds to human perceptual judgments, thus implying better visual appearance.
@@
NeRF [16] (ACM 21) $26.50$ $25.17$ $27.40$ $31.16$ $27.45$ $20.92$ $20.36$ $32.70$ $26.80$
TensoRF [6] (ECCV’22) $26.73$ $25.27$ $\bm{28.60}$ $31.36$ $28.14$ $21.30$ $19.87$ $32.35$ $26.97$
Plenoxels [8] (CVPR’22) $26.29$ $25.46$ $27.83$ $31.09$ $27.58$ $21.41$ $20.24$ $30.22$ $26.48$
Mip-NeRF 360 [4] (CVPR’22) -- $24.59$ $27.56$ $31.34$ $\bm{28.51}$ $19.84$ $19.51$ $\bm{33.49}$ $\bm{27.86}$
L2G-NeRF [7] (CVPR’23) $24.54$ $24.57$ $24.90$ $29.27$ $23.12$ $19.02$ $19.71$ $32.25$ $23.49$
BioNeRF (Ours) $\bm{27.01}$ $\bm{26.51}$ $27.89$ $\bm{32.34}$ $27.99$ $\bm{22.23}$ $\bm{20.80}$ $30.75$ $27.56$
SSIM $\uparrow$
@@
NeRF [16] (ACM 21) $0.811$ $0.792$ $0.827$ $0.881$ $0.828$ $0.690$ $0.641$ $0.948$ $0.880$
TensoRF [6] (ECCV’22) $0.839$ $0.814$ $0.871$ $0.897$ $0.877$ $0.752$ $0.649$ $0.952$ $0.900$
Plenoxels [8] (CVPR’22) $0.839$ $0.832$ $0.862$ $0.885$ $0.857$ $0.760$ $0.687$ $0.937$ $0.890$
Mip-NeRF 360 [4] (CVPR’22) -- $0.820$ $0.867$ $0.900$ $\bm{0.909}$ $0.721$ $0.660$ $\bm{0.965}$ $\bm{0.927}$
L2G-NeRF [7] (CVPR’23) $0.750$ $0.750$ $0.740$ $0.840$ $0.740$ $0.560$ $0.610$ $0.950$ $0.800$
BioNeRF (Ours) $\bm{0.861}$ $\bm{0.837}$ $\bm{0.873}$ $\bm{0.914}$ $0.882$ $\bm{0.796}$ $\bm{0.714}$ $0.956$ $0.911$
LPIPS $\downarrow$
@@
NeRF [16] (ACM 21) $0.250$ $0.280$ $0.219$ $0.171$ $0.268$ $0.316$ $0.321$ $0.178$ $0.249$
TensoRF [6] (ECCV’22) $0.124$ $0.155$ $0.106$ $0.075$ $0.123$ $0.153$ $0.201$ $0.082$ $0.099$
Plenoxels [8] (CVPR’22) $0.210$ $0.224$ $0.179$ $0.180$ $0.231$ $0.198$ $0.242$ $0.192$ $0.238$
Mip-NeRF 360 [4] (CVPR’22) -- $0.210$ $0.140$ $0.117$ $0.124$ $0.231$ $0.246$ $0.115$ $0.158$
L2G-NeRF [7] (CVPR’23) $0.200$ $0.260$ $0.170$ $0.110$ $0.260$ $0.330$ $0.250$ $0.080$ $0.160$
@@
5.3 Qualitative evaluation.
This section evaluates the quality of the views generated by the proposed method in terms of visual aspects. Figure 2 compares BioNeRF against three baselines, i.e., NeRF [16], Mip-NeRF 360 [4], and TensoRF [6] [^5] Baseline views obtained from https://nerfbaselines.github.io/. considering a Blender dataset’s Lego view. Compared to NeRF, BioNeRF presents a result closer to the ground truth, especially if the toy’s grid and piston reflections are analyzed in the bottom image. Mip-NeRF 360 shows some black artifacts in the white area on the top image. Regarding TensoRF, BioNeRF shows itself more capable of extracting the reflexible components and reproducing the floor color with better precision.
Lego
NeRF
MipNeRF 360
TensoRF
BioNeRF
GT
Figure 2: Qualitative results of BioNeRF and comparison methods, namely NeRF [16], Mip-NeRF 360 [4], and TensoRF [6], as well as the ground truth images (GT) on two Blender dataset’s scenes.
Regarding the LLFF dataset, Figure 3 compares BioNeRF against Mip-NeRF 360 [4] and TensoRF [6] considering a Fern view. In This context, BioNeRF views look cleaner than the instances generated by TensoRF, whose images present some high-frequency artifacts. Regarding Mip-NeRF 360, although the images’ resolution looks a bit better than BioNeRF, they show some distorted details considering the air grid on the bottom picture.
Fern
MipNeRF 360
TensoRF
BioNeRF
GT
Figure 3: Qualitative results of BioNeRF and comparison methods, namely, Mip-NeRF 360 [4] and TensoRF [6], as well as the ground truth images (GT) on LLFF dataset’s Fern scenes.
5.4 Ablation study
This section provides an evaluation of the BioNeRF effectiveness over the Blender dataset considering three implemented versions, based on the standard NeRF [16], NeRFacto [^6] https://docs.nerf.studio/nerfology/methods/nerfacto.html, and TensoRF [6], which will be referred here as BioNeRF, BioNeRFacto, and BioTensoRF, respectively. The biologically plausible module is implemented in a substructure of NeRF’s pipeline called field. In this context, BioNeRF changes are implemented in the Coarse and Fine blocks of NeRF’s pipeline [^7] https://docs.nerf.studio/nerfology/methods/nerf.html. Considering BioNeRFacto and BioTensoRF, such changes are implemented in NeRFacto’s nerfacto field and TensoRF’s field, respectively.
Table 4 provides the average results achieved after training the models for $50k$ iteration. BioNeRF and BioTensoRF outperformed BioNeRFacto results, considering all the evaluation metrics. Considering both methods, BioNeRF outperformed BioTensoRF by $0.001$ considering the SSIM and a far advantage considering the LPIPS metric, While BioTensoRF obtained a slightly better PSNR result. Running the training during more iterations for some dataset scenes, we could observe that BioNeRF sometimes surpasses BioTensoRF PSNR values.
Table 4: Average results obtained considering three distinct implementations of BioNeRF over Blender datasets.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
BioNeRFacto $19.01$ $0.792$ $0.216$
BioTensoRF $29.15$ $0.928$ $0.063$
BioNeRF $28.81$ $0.929$ $0.031$
@@ 6 Discussions, Conclusions, and Future Works
Future Works. Regarding future works, we aim to improve the model by introducing more elements from a biologically plausible perspective, like spiking neurons, implemented in SpikingNeRF [25]. Additionally, we strive to extend our model to infer in real time and learn from a few shots.